3. Evaluate S (2 + 2)dz, where C is the line segment from 0 to 1+i.
2020:1 Spring, MATH5880:001 Complex Variables
Answer #1

- I - J (2 + 2 ) dz 2= xriy z = x-diy. (2+2 = dx ] I = (dx dr , dz= carady. I= Sax (antudy) co to Hi = along & y=x line (oo)